2) Decomat Ltd (D) acquires 75% of equity shares in Macs Ltd (M) through exchanging one $1 share in D for three $1 shares in M. M has a total of 60 000 shares. D’s shares have a market value of $3 on the date of acquisition, while M’s shares are valued at $2. Determine the cost of the business combination. A) $15 000 B) $45 000 C) $60 000 D) $90 000
Answer #1

The correct answer for this is B) $45000

See the below calculation

Decomat Acquired 75% of share in Macs 45000 shares

Decomate one share market price is $3

Macs one share market value is $2

Decomate exchanges 1 share for 3 Macs share (Decomate15000 share for 45000 share of Macs)

Decomate spent 15000 * 3 = 45000

But now the value off Macs 45000 share is 45000 * 2 = 90000

So 90000 - 45000 = 45000